OpenStudy (anonymous):

A goes through 2 3/5 boxes of nails in 3 1/2 rooves. How much would he use to finish 8 rooves?

OpenStudy (mathstudent55):

By "rooves", do you mean "roofs", the plural of "roof"?

OpenStudy (mathstudent55):

Use a proportion. Let x = the unknown number of boxes needed for 8 roofs. \(\large 2\dfrac{3}{5}\) boxes are to \(\large 3 \dfrac{1}{2}\) roofs as x boxes are to 8 roofs. \(\Large \dfrac{2\frac{3}{5}}{3 \frac{1}{2}} = \dfrac{x}{8}\) Convert the mixed numerals into fractions. Then cross multiply and solve for x.

OpenStudy (knightofpentacles):